OpenStudy (uri):

How many different committies of 3 men and 4 girls can be formed from 8 men and 6 girls.

OpenStudy (shamim):

\[C _{3}^{8} \times C _{4}^{6}\]

OpenStudy (shamim):

C= combination

OpenStudy (uri):

Have you written the formula the way it should be? @shamim

terenzreignz (terenzreignz):

ohh combinations this time :D You have an idea what @shamim meant? :D This is \[\Large ^nC_r\] this time

OpenStudy (shamim):

ya

OpenStudy (uri):

Yes! Got it.

terenzreignz (terenzreignz):

Okay, admire this formula this time around :D \[\huge ^nC_r = \frac{n!}{(n-r)!\color{red}{r!}}\]

terenzreignz (terenzreignz):

I put the r! at the bottom in red to emphasise its difference from nPr

OpenStudy (shamim):

ya i was meaning it

terenzreignz (terenzreignz):

You DO remember what factorial means, right @uri ? :)

OpenStudy (uri):

Yes i do remember :)

terenzreignz (terenzreignz):

Good. That's all I want to know :P

OpenStudy (uri):

But tell me how will they be written in the formula @terenzreignz

OpenStudy (uri):

@msingh Can you please give me the link again? :3

terenzreignz (terenzreignz):

Let me explain the difference between nPr and nCr in a 'nice' manner. Say you have 6 numbers 1 2 3 4 5 6 And you're asked how many THREE DIGIT NUMBERS you can form using these 6 numbers. The answer would be nPr examples would be 123 124 126 231 etc

OpenStudy (uri):

Hey amistre explained me the difference yesterday :D @terenzreignz I know

terenzreignz (terenzreignz):

all right, fine :/

terenzreignz (terenzreignz):

then for the first one \[\Large ^8C_3 = \frac{8!}{(8-3)!3!}= \frac{8!}{5!\cdot 3!}\]

OpenStudy (cwrw238):

one way to work out combinations ( if your calculator doesn't help) is as follows 8C3 = 8 * 7 * 6 ------- = 56 3 * 2 * 1

OpenStudy (cwrw238):

8C2 = 8*7 --- = 28 2* 1

OpenStudy (uri):

7C3 =7P3/3 =210/3! = :35 Just an example,We have to do it like that? @terenzreignz

terenzreignz (terenzreignz):

Me and my formulas :D Looks like @cwrw238 is showing us the simpler way to do it :) if 8P3 has us picking three numbers, downwards (to 0) starting from 8... 8P3 = 8 x 7 x 6 8C3 would have a fraction, the numerator having us pick three numbers downwards starting from 8... while the denominator has us pick three numbers *upwards starting from 1* Like so: \[\huge ^8C_3=\frac{8\cdot7\cdot6}{1\cdot2\cdot3}\]

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

also its worth knowing that 8C5 = 8C(8-5) = 8C3

OpenStudy (cwrw238):

7C6 = 7C1 = 7

OpenStudy (uri):

8C3 =8P3/3! =336/3! =56 right? :D

OpenStudy (cwrw238):

yes

terenzreignz (terenzreignz):

Yeah, on the whole, yes :D \[\huge ^nC_r = \frac{^nP_r}{r!}\]

OpenStudy (uri):

And for the other one : 6C4 =6P4/4! =360/4!...right? But before that I have a Question....3! =6 right? so what does 4! mean?

OpenStudy (cwrw238):

4! = 4*3*2*1

OpenStudy (uri):

Got it :D

terenzreignz (terenzreignz):

4! = 4 times 3! haha That applies to all positive integers by the way... \[\huge n! = n\times (n-1)!\]

OpenStudy (uri):

6C4 =6P4/4! =360/4! =15..?

terenzreignz (terenzreignz):

It's like slowly unraveling a tape or a roll of thread... \[\large 8!=8\cdot 7!\]\[\large 8!=8\cdot 7\cdot6!\]\[\large 8!=8\cdot 7\cdot6\cdot5!\]\[\large 8!=8\cdot 7\cdot6\cdot5\cdot4!\]

terenzreignz (terenzreignz):

That is correct :)

OpenStudy (uri):

OMG YAAAY @terenzreignz xD

terenzreignz (terenzreignz):

:) Kudos @shamim @cwrw238 ^.^

OpenStudy (uri):

Terence give @cwrw238 a medal foh meh :D

OpenStudy (cwrw238):

lol ty

OpenStudy (uri):

:3